OpenStudy (anonymous):

solve for x: -3(x+3)=-3(x+1)-5

OpenStudy (kamibug):

1. Distribute the -3 on both sides. -3x-9 = -3x-3-5 2. Combine Like Terms on the right. -3x-9 = -3x-8 3. Isolate the variable. -3x-9 = -3x-8 -3x-1 = -3x -1=0 No solution.